OpenStudy (anonymous):

Please Help! Find all real solutions of the equation. 1/x-1-2/x^2=0

OpenStudy (abb0t):

Here's the start: \[\frac{ x^2-2(x-1) }{x^2(x-1)}=0\]

OpenStudy (anonymous):

so then would it be \[x^2-2x+1=0 \]

OpenStudy (anonymous):

-2(x-1) = -2x + 2

OpenStudy (anonymous):

How did u get that stiwan

OpenStudy (anonymous):

hello

OpenStudy (abb0t):

No, you are correct @x0hill. Distribute the 2 across the parenthesis now.

OpenStudy (anonymous):

what do you mean distribute the 2

OpenStudy (abb0t):

Well, you have: \[x^2-2(x-1) = x^2... = 0\]

OpenStudy (abb0t):